# Not All Attention Heads Are What You Need: Refining CLIP’s Image Representation with Attention Ablation

### v1

@@ Abstract
This paper studies the role of attention heads in CLIP’s image encoder. While CLIP has exhibited robust performance across diverse applications, we hypothesize that certain attention heads negatively affect final representations and that ablating them can improve performance in downstream tasks. To capitalize on this insight, we propose a simple yet effective method, called Attention Ablation Technique (AAT), to suppress the contribution of specific heads by manipulating attention weights. By integrating two alternative strategies tailored for different application scenarios, AAT systematically identifies and ablates detrimental attention heads to enhance representation quality. Experiments demonstrate that AAT consistently improves downstream task performance across various domains, boosting recall rate by up to 11.1% on CLIP-family models for cross-modal retrieval. The results highlight the potential of AAT to effectively refine large-scale vision-language models with virtually no increase in inference cost.
@@ 1 Introduction
As a pioneering large-scale vision-language model (VLM), CLIP [28] has garnered widespread attention for its simple yet effective design and impressive performance across a wide range of downstream tasks [16, 40, 25, 6]. While early research focused on improving CLIP through advancements in data [20], supervision [38], and architecture [16], recent studies have shifted toward analyzing its learned representations [11, 18, 1]. These efforts aim to uncover the intrinsic characteristics of CLIP’s representations, providing insights for further enhancement.
A recent study [11] investigates CLIP’s visual representations by decomposing them across attention heads, revealing a strong correlation between certain attention heads and specific visual concepts. By manually removing heads associated with spurious cues in the last four transformer layers, the CLIP’s visual representation is improved for a targeted zero-shot classification task [11]. This suggests that individual attention heads contribute uniquely to the output representation, act as property-specific “filters” for visual concepts. Building on this insight, we systematically extend the analysis to all attention heads in CLIP’s image encoder across diverse downstream tasks.
Trained on vast amounts of internet-sourced image-text pairs, the CLIP-family models may include attention heads that encode task-irrelevant signals—such as domain biases or spurious cues—that hinder downstream performance. Additionally, some heads might overfit to noise inherent in the uncurated training data. These detrimental heads can be distributed across different transformer layers in the image encoder. We hypothesize that ablating such heads could refine CLIP’s representations, thereby consistently improving performance across diverse downstream tasks.
To verify our hypothesis, we conduct preliminary experiments assessing the impact of individual attention heads. We find that ablating even a single detrimental head can enhance CLIP’s downstream performance, with further gains from ablating simple combinations of negatively impactful heads. Inspired by this, we develop an efficient Attention Ablation Technique (AAT) that refines CLIP representations by evicting groups of detrimental heads. AAT formulates the problem as a combinatorial optimization task, solved using either a Genetic Algorithm (GA) or Back-Propagation (BP).
The key difference between GA-based AAT (AAT-GA) and BP-based AAT (AAT-BP) lies in how they identify detrimental attention heads. AAT-GA uses a fitness function that measures the distance between positive image-text pairs and hard negatives, while AAT-BP introduces a lightweight training scheme with a learnable gating parameter for each head to determine its relevance. Both strategies yield remarkable improvements in multi-domain retrieval and classification tasks, with their distinct advantages making them suitable for different use cases.
AAT leaves CLIP’s model parameters unchanged, instead manipulating attention weights of selected heads, as shown in Figure 1. It suppresses image token weights to a low level while amplifying the class token weight as compensation. This ensures minimal image token contribution from ablated heads to the output representation. As an agile and versatile technique, AAT offers potential advantages in resource-constrained scenarios, such as limited compute or scarce high-quality data, while preserves strong generalization across diverse domains. Our main contributions are as follows:
  • We extensively explore the roles of attention heads in CLIP’s image encoder and show that ablating detrimental heads improves representation quality. To our knowledge, this is the first systematic study on refining CLIP’s outputs by probing attention head impacts.
  • We introduce AAT, a simple yet effective method for automatically identifying and ablating detrimental attention heads without altering model parameters. AAT offers two alternative strategies—GA and BP—tailored for different application scenarios.
  • AAT achieves consistent and significant improvements in diverse downstream tasks, delivering recall rate gains of up to 11.1% for retrieval, with little additional inference cost.
@@
Figure 1: An illustration of AAT-improved CLIP for text-to-image retrieval. “A-head $i$” denotes the $i$-th head in MHA. With model parameters frozen, AAT suppresses the selected detrimental heads in the image encoder. Cross marks indicate ablated heads while check marks denote retained ones.
@@ 2 Related Work
Recent studies explore CLIP’s representations from various perspectives. [26] analyzes entanglement between words and images. [31] finds a strong connection between the modality gap and loss local minima. [22] reports a semantic shift toward background regions. [39] improves image-text matching explainability via gradients and heatmaps. [7] finds that CLIP exhibits distinctive outlier features tied to ImageNet shift robustness. Closer to AAT, [3] boosts interpretability by converting dense embeddings into sparse, concept-based ones, and [11] reveals spatial localization and property-specific roles of attention heads.
@@
The attention mechanism is central to transformer-based models [4], and manipulating it in pre-trained models has shown notable benefits. [37] calibrates weights toward hidden attention sinks in LLMs; [18] reformulates attention weights for segmentation; [17] reweights semantic tokens to refine CLIP’s text embeddings; and [24] amplifies image token weights to alleviate hallucinations in VLMs. AAT shares conceptual similarities with these heuristic methods but introduces a novel manipulation strategy grounded in systematic empirical exploration.
Attention Head Pruning
The idea behind AAT bears some resemblance to early studies on language models (LMs) [27, 33, 2], which identify redundancy among attention heads. These methods use pruning to expedite inference, typically at the cost of slight quality loss. While head pruning has been extensively explored in LMs, its exploration in VLMs remains limited.
@@
PEFT [14, 19, 13, 32] are widely used in resource-constrained scenarios by updating only a small subset of model parameters. The proposed AAT-BP shares some similarities with PEFT, as it introduces hundreds of learnable parameters. However, key differences remain: PEFT typically functions as a black box, with the role of each parameter unclear, whereas AAT-BP offers explicit interpretability of every parameter. Moreover, AAT-BP is more parameter-efficient, requiring orders of magnitude fewer parameters than typical PEFT approaches [9].
3 The Impact of Individual Attention Heads
In a preliminary experiment, we probe the impact of each individual attention head in CLIP’s image encoder. Employing the technique described in Section 4.1, we ablate one head at a time in a CLIP model and assess its performance on a retrieval task. For instance, in ViT-B-based CLIP with 144 heads across 12 layers, this grid search yields 144 results—some outperforming the original model, others not. The top 5 highest-performing configurations are ranked in Table 1, with full results in Appendix B. Notably, ablating even a single detrimental head can boost mean recall by up to 1.2%.
Building on this observation, we conduct an experiment that ablates a straightforward combination of all individually identified detrimental heads—those that improve recall when ablated alone. As shown in Table 2, this naive combination consistently boosts retrieval performance across multiple test subsets. Interestingly, although the heads are selected independently, their joint ablation yields further gains. This motivates the development of more effective strategies for optimal head selection.
@@ 3 The Impact of Individual Attention Heads
Table 1: Top 5 single-head ablation configurations for text-to-image retrieval on the COCO-CN $val$ set [21]. “m-n” demotes the $n$-th head in the $m$-th transformer layer. Evaluations are based on a ViT-B-based CLIP model [35], using mean-R (the average of R@1, R@5, and R@10) as the metric.
Models $val$ $test$ $all$
vanilla CLIP [35] 79.9 81.1 41.7
@@
Table 2: Comparison of mean-R for text-to-image retrieval on the COCO-CN $val$, $test$ and $all$ set.
@@ 4 Attention Ablation Technique
Figure 2: The left $5\times 5$ matrix represents the original attention weight matrix, while the right one shows the attention weights after manipulating. CLS and IMG denote the class token and an image token, respectively, with the input token sequence length being 5. Deeper blue indicates lower attention scores approaching zero, while deeper red represents higher scores closer to one.
4.1 Ablating the Attention Head
A transformer layer consists of a multi-head attention module (MHA) and a feed-forward network (FFN) [10]. An input token sequence of length $N$ is structured with a leading class token followed by $N-1$ image tokens. Let $x^{l}_{h}$ represent the input to the $h$-th attention head of the MHA in the $l$-th transformer layer. The output of each head in the MHA ${y}^{l}_{h}$ can be formulated as follows:
@@ 4.1 Ablating the Attention Head
where $f_{q}$, $f_{k}$, and $f_{v}$ denote the projection layers for queries, keys, and values, respectively. $LN$ refers to LayerNorm, $d_{k}$ is the embedding dimension of each head, and $\mathcal{A}^{l}_{h}$ denotes the attention weight matrix with dimensions $N\times N$.
As shown in the formulations above, the attention weight matrix $\mathcal{A}$ is generated by a softmax function, ensuring that each row sums to 1. The output of each attention head, ${y}$, is computed as the matrix product of $\mathcal{A}$ and the projected values. To ablate an head, we manipulate $\mathcal{A}$ to suppress the contribution of image tokens in the output representation. An illustrative example is provided in Figure 2. Specifically, this manipulation involves the following two steps:
@@
    Reduce the importance of image tokens: Adjust $\mathcal{A}$ as $\mathcal{A}[:,j]$leftarrow $\mathcal{A}[:,j]\times\beta$, where $j\in\{1,2,...,N-1\}$.
@@
    Reweight each row of $\mathcal{A}$: Normalize $\mathcal{A}$ by setting $\mathcal{A}[i,j]$leftarrow $\mathcal{A}[i,j]~{}/~{}\mathcal{A}_{i}$, where $\mathcal{A}_{i}=\sum_{j}\mathcal{A}[i,j]$ and $i,j\in\{0,1,...,N-1\}$. This ensures that each row still sums to 1.
@@
We use a hyper-parameter $\beta$ to control the degree of suppression applied to the weights of image tokens. Unless otherwise noted, we set $\beta=0.1$. After applying the two manipulation steps, the manipulated attention weight matrix replaces the original, while all other operations in MHA remain unchanged. This enables refinement of the output representation without altering model parameters.
4.2 AAT with Genetic Algorithm
As discussed in Section 3, ablating detrimental attention heads in CLIP can refine its output representation. Our objective is to identify an optimal selection of such heads. We formulate this as a combinatorial optimization problem and address it using a small validation set $\mathcal{D}$ containing a limited number of image-text pairs. We use a Genetic Algorithm (GA) [12], which evolves candidate solutions over generations to navigate the complex search space and acquire high-quality solutions.
Specifically, we represent all attention heads in the image encoder using a binary vector, where each element corresponds to a head (1 for ablated, 0 for retained). Using the validation set $\mathcal{D}$, GA is applied to optimize this vector configuration. To align with CLIP’s training objective that brings matching image-text pairs closer while pushes apart mismatched ones, we design a fitness function $\mathcal{F}$ that reflects this property. $\mathcal{F}$ serves as the optimization target for GA and is defined as follows:
where $N$ is the number of image-text pairs in $\mathcal{D}$. $S_{pos}^{i}$ is cosine similarity between the $i$-th text and its ground truth image. $S_{neg}^{i,j}$ is cosine similarity between the $i$-th text and the $j$-th image from $\mathcal{H}^{i}$, an updatable set of hard negatives for the $i$-th text. We detail the construction of $\mathcal{H}^{i}$ below.
Initially, we evaluate the vanilla CLIP on $\mathcal{D}$ for text-to-image retrieval. For each text query $i$, the top $k_{1}$ non-matching images (i.e., false positives) are collected as hard negatives to form the basis of $\mathcal{H}^{i}$. To avoid overfitting to these cases, we augment $\mathcal{H}^{i}$ with $k_{2}$ randomly sampled non-matching images from $\mathcal{D}$. At the end of each GA generation, these $k_{2}$ random samples are refreshed with newly sampled ones. This dynamic $\mathcal{H}$ is crucial to AAT-GA, as it encourages larger distance margins between true matches and evolving hard negatives, while also considering newly emerging challenging cases.
In Eq. 3, the first term of $\mathcal{F}$ promotes similarity between positive image-text pairs, while the second penalizes similarity with the hardest negatives in $\mathcal{H}$. Together with standard crossover and mutation operations, $\mathcal{F}$ effectively guides GA toward an optimal or near-optimal head selection for ablation.
Application Scenarios
AAT-GA refines CLIP’s representations entirely at inference time, making it well-suited for resource-constrained scenarios. It can be deployed on low-power, inference-targeted edge devices with limited-precision compute capabilities, such as INT8/INT16 chips (e.g., Google Coral Edge TPU, Ethos-N78, Mythic M1108 AMP). Moreover, AAT-GA is effective in data-scarce scenarios, requiring only a small number of samples for optimization (see Section 5.3.2), whereas supervised finetuning (SFT) tends to overfit under such conditions (see Appendix F).
4.3 AAT with Back-Propagation
Since GA may require more inference trials with a poorly initialized population, and not all resource-constrained scenarios lack floating-point (FP) computation power, we propose an alternative strategy for identifying detrimental heads using gradients: Back-Propagation (BP) optimization. Unlike the uniform suppression factor $\beta$ used in AAT-GA, BP introduces an individual learnable parameter $\alpha_{i}$ for each probed head. We replace $\beta$ with a head-specific factor $\beta_{i}\in[0,1]$, which determines the degree to which the $i$-th head is ablated. $\beta_{i}$ is defined as follows:
$\beta_{i}$ $=\text{Sigmoid}(\tau\cdot\alpha_{i})$ (4)
where the temperature $\tau$ is set above 1 to sharpen the sigmoid output, encouraging $\beta_{i}$ to approach extreme values (0 or 1) during training. Empirically, we set $\tau=5.0$ and initialize each $\alpha_{i}$ to 1.0, yielding an initial $\beta_{i}$ close to 1. As in AAT-GA, we use the same validation set $\mathcal{D}$ to optimize the parameters ${\alpha_{i}}$, guided by the contrastive supervision objective used in CLIP.
Application Scenarios
While BP might get stuck in a local optimum, its fast convergence makes it a practical and efficient solution for real-world applications. Its low compute and data requirements enhance its applicability for resource-limited settings. In spirit, AAT-BP resembles a PEFT method, yet it includes several orders of magnitude fewer learnable parameters than typical PEFTs [14, 13]. We believe this extreme efficiency stems from a deep understanding of CLIP’s attention heads.
@@ 4.3 AAT with Back-Propagation
Table 3: Comparison on MS COCO for retrieval. R@k denotes recall rate at rank k, and mR (namely mean-R) denotes the average of R@1, R@5, and R@10. Model sizes are denoted as B (ViT-B), L (ViT-L), and H (ViT-H). Best results are highlighted in bold. No further elaboration in the following.
5 Experiments
5.1 Experimental Settings & Implementation Details
Models
We evaluate AAT on CLIP [28] and Chinese-CLIP [35], across ViT-B, ViT-L, and ViT-H. Specifically, we use OpenCLIP released models [15] trained on LAION-2B (a subset of LAION-5B [30]) and Chinese-CLIP released models trained on about 200M Chinese image-text pairs.
AAT Optimization Data
AAT-GA Implementation
The key hyper-parameters for GA are as follows: the population size is set based on the number of attention heads—48 for ViT-B, 96 for ViT-L, and 128 for ViT-H. Two-point crossover is applied with a probability of 0.9, and mutation occurs with a probability of 0.5 using a flip-bit strategy. Tournament selection with a size of 3 is used. The evolution runs for up to 100 generations, with early stopping triggered by stagnant fitness gains or low population diversity.
AAT-BP Implementation
The settings for training $\alpha$ in BP are as follows: a learning rate of 2e-2 for CLIP and 5e-2 for Chinese-CLIP, with a batch size of 256. ViT-B-based models are trained for 32 epochs, and larger models for 64. Other hyper-parameters follow the OpenCLIP codebase [15].
Benchmarks
We benchmark AAT-improved CLIP on MS COCO [23], Flickr30k [36], and ReCoS [5], and evaluate AAT-improved Chinese-CLIP on two widely used Chinese retrieval datasets: COCO-CN [21] and Flickr30k-CNA [34]. For MS COCO, Flickr30k, COCO-CN, and Flickr30k-CNA, we create two evaluation splits: the test set (i.e., the original test split) and the all set (including all samples in the dataset). This setup enables a comprehensive analysis of AAT in large-scale retrieval. To further broaden the evaluation, we also assess AAT on the val set of ImageNet-1k [8] for zero-shot classification. Details on datasets and evaluation metrics are provided in Appendix A.

### v2

@@ Abstract
This paper investigates the role of attention heads in CLIP’s image encoder. Building on interpretability studies, we conduct an exhaustive analysis and find that certain heads, distributed across layers, are detrimental to the resulting representations. To mitigate their impact, we propose a simple yet effective Attention Ablation Technique (AAT) that suppresses selected heads by directly manipulating their attention weights. By incorporating two complementary strategies tailored to different application scenarios, AAT enables the systematic identification and ablation of harmful heads with minimal overhead. Experiments show that AAT consistently improves downstream performance across diverse domains, boosting recall by up to 11.1% on cross-modal retrieval benchmarks. These results highlight that AAT can effectively refine large-scale VLMs with virtually no extra inference cost, while yielding semantically meaningful patterns that align with existing interpretability findings.
@@ 1 Introduction
As a pioneering large-scale vision-language model (VLM), CLIP [radford2021learning] has garnered widespread attention for its simple yet effective design and impressive capability across a wide range of downstream tasks [jia2021scaling, zhou2022conditional, luddecke2022image, cho2021unifying]. Early CLIP research focuses on advances in data [li2021align], supervision [zhai2023sigmoid], and architecture [jia2021scaling], recent studies have shifted toward analyzing its learned representations [gandelsmaninterpreting, lan2025clearclip, abbasi2025deciphering]. These efforts aim to uncover the intrinsic characteristics of CLIP’s representations, providing insights for further enhancement.
@@
Figure 1: An illustration of AAT-improved CLIP for text-to-image retrieval. “A-head ii” denotes the $i$-th head in MHA. With model weights frozen, AAT ablates the selected image encoder’s heads. Cross marks denote head ablation while check marks for retention.
A recent interpretability study [gandelsmaninterpreting] investigates CLIP’s image representations by decomposing them across attention heads, revealing strong correlations between specific heads and particular linguistic concepts. Through manually removing those heads associated with spurious cues in the last four transformer layers, it improves CLIP’s image representations for a targeted zero-shot classification task. This finding suggests that individual head makes distinct contribution to output representations, acting as property-specific “filters” over visual concepts. Building on this insight, we systematically extend the analysis to all attention heads in CLIP’s image encoder across diverse downstream tasks.
Trained on vast amounts of internet-sourced image–text pairs, CLIP models probably develop attention heads that encode task-irrelevant signals—such as domain biases or spurious cues—that may hinder downstream performance. Some heads may also overfit to noise inherent in uncurated training data. Such detrimental heads can appear throughout the transformer layers of the image encoder. We hypothesize that ablating them can refine the output and thereby yield consistent gains across diverse downstream tasks.
To verify our hypothesis, we conduct a preliminary experiment to assess the impact of individual attention heads. We find that ablating even a single detrimental head can enhance CLIP’s downstream performance, with further gains from ablating simple combinations of negatively impactful heads. Inspired by this, we develop an efficient Attention Ablation Technique (AAT) that refines the representations by evicting groups of detrimental heads. It formulates the problem as a combinatorial optimization task, solved using either Genetic Algorithm (GA) or Back-Propagation (BP).
The key difference between GA-based AAT (AAT-GA) and BP-based AAT (AAT-BP) lies in how they identify detrimental attention heads. AAT-GA uses a fitness function that measures the distance between positive image–text pairs and hard negative pairs, whereas AAT-BP introduces a lightweight training scheme with a learnable gating parameter for each head to assess its relevance. Both strategies yield substantial gains on multi-domain retrieval and classification tasks, with their respective strengths making them suitable for different deployment scenarios.
AAT leaves CLIP’s parameters unchanged and instead operates by manipulating the attention weights of selected heads, as illustrated in Figure 1. It suppresses the weights on image tokens while correspondingly amplifying the class token weight, ensuring that ablated heads contribute minimally to the output representation. As an agile and versatile technique, AAT is particularly well suited to resource-constrained scenarios, like limited compute or scarce high-quality data, while preserving strong generalization across diverse domains. The main contributions as follows:
  • We extensively explore the impacts of attention heads in CLIP’s image encoder and show that ablating detrimental heads improves representation quality. To our knowledge, this is the first systematic study of refining CLIP’s outputs by probing the impact of individual attention heads.
  • We introduce AAT, a simple yet effective method for automatically identifying and ablating detrimental attention heads without altering original model weights, offering two variants tailored to different application scenarios.
  • AAT delivers consistent and substantial gains across diverse downstream tasks, improving recall by up to 11.1% on retrieval benchmarks with negligible extra overhead.
  • Out result analysis provides concrete evidence that attention heads are semantically meaningful, reinforcing and extending existing interpretability findings of VLMs.
@@ 2 Related Work
Recent studies investigate CLIP’s representations from various perspectives. [materzynska2022disentangling] analyzes entanglement between words and images. [li2023clip] presents a semantic shift toward background regions. [shi2023towards] finds a strong connection between the modality gap and the loss local minima. [zhao2024gradient] improves image-text matching explainability via gradients and heatmaps. [crabbeinterpreting] finds that CLIP exhibits distinctive outlier features tied to ImageNet shift robustness. Most closely related to AAT, [bhalla2024interpreting] boosts interpretability by converting dense embeddings into sparse, concept-based ones, and [gandelsmaninterpreting] reveals spatial localization and property-specific roles of attention heads.
@@
Attention mechanism is central to Transformers [brauwers2021general], and manipulating attention weights in trained models has shown notable benefits. Prior heuristic method calibrates its weights toward hidden attention sinks in LLMs [yu2024unveiling]; reformulates attention for segmentation [lan2025clearclip]; reweights semantic tokens to refine text embeddings [kim2024semantic]; and amplifies image-token weights to mitigate hallucinations in VLMs [liu2025paying]. AAT is conceptually related to them but introduces a novel, empirically grounded manipulation strategy based on attention-head behaviors.
Attention Head Pruning in LMs
The idea behind AAT parallels early works on language models [michel2019sixteen, voita2019analyzing, behnke2020losing], which reveal redundancy among attention heads. Those methods typically prune heads to slim models with modest accuracy loss. Though head pruning is well established for LMs, its study to much compact VLMs remains underexplored.
@@
AAT-BP is superficially similar with PEFT [hulora, lester2021power, houlsby2019parameter, sung2022lst, wu2024reft]—requiring only hundreds of learnable gates—but, unlike PEFT’s largely black-box updates, each parameter in AAT-BP is explicitly interpretable. It is also far more parameter-efficient, requiring orders of magnitude fewer parameters than PEFT [ding2023parameter].
@@
Table 1: Top five single-head ablation configurations for text-to-image retrieval on COCO-CN valval set [li2019coco]. “m-n” demotes the $n$-th head in the $m$-th layer. Evaluations are based on CLIP-ViT-B [yang2022chinese], mean-R (the mean of R@1, R@5, and R@10) as metric.
Models valval testtest allall
vanilla CLIP [yang2022chinese] 79.9 81.1 41.7
@@
Table 2: Comparison of mean-R for text-to-image retrieval on the COCO-CN valval, testtest and allall set.
3 The Impact of Individual Attention Heads
As a preliminary step, we conduct a simple head-wise ablation sweep of CLIP’s image encoder. Using AAT (described in Section 4.1), we ablate one attention head at a time and evaluate cross-modal retrieval performance. For ViT-B with 144 heads across 12 layers, this yields 144 results, several of which outperform the original baseline. The top 5 highest-performing configurations are ranked in Table 1, with full results in Appendix 8. Remarkably, ablating even a single detrimental head can raise mean recall by up to 1.2%.
Building on this observation, we ablate the direct union of all the individually identified detrimental heads—those whose single-head ablation improves recall. As shown in Table 2, this naive combination consistently boosts retrieval performance across multiple test subsets. Notably, although these heads are selected independently, their joint ablation yields further gains. This motivates us to develop more principled strategies for optimal head selection.
@@ 4 Attention Ablation Technique
4.1 Ablating a Single Attention Head
A transformer layer consists of a multi-head attention module (MHA) and a feed-forward network (FFN) [dosovitskiy2020image]. For the CLIP’s image encoder, the input token sequence of length $N$ is composed of a leading class token followed by N1N-1 image tokens. Let $x^{l}_{h}$ represent the input to the $h$-th attention head of MHA in the $l$-th transformer layer. The output of each head, ${y}^{l}_{h}$, can be formulated as follows:
@@ 4.1 Ablating a Single Attention Head
where $f_{q}$, $f_{k}$, and $f_{v}$ denote projection layers for queries, keys, and values, respectively. $LN$ refers to LayerNorm, $d_{k}$ is the embedding dimension of each head, and $\mathcal{A}^{l}_{h}$ denotes the attention weight matrix with dimensions $N\times N$.
As shown in the formulations above, the attention weight matrix $\mathcal{A}$ is produced by a softmax function, ensuring that each row sums to 1. The output of each attention head, y{y}, is then obtained by multiplying $\mathcal{A}$ with the projected values. To ablate a specific head, we manipulate $\mathcal{A}$ to suppress the contribution of image tokens to the output representation, as illustrated in Figure 2. Specifically, this involves two steps:
@@
    Reducing the importance of image tokens: Modify $\mathcal{A}$ as $\mathcal{A}[:,j]\leftarrow\mathcal{A}[:,j]\times\beta$, where $j\in\{1,2,...,N-1\}$.
@@
    Renormalizing each row of $\mathcal{A}$: Scale $\mathcal{A}$ as $\mathcal{A}[i,j]\leftarrow\mathcal{A}[i,j]/\mathcal{A}_{i}$, where $i,j\in\{0,1,...,N-1\}$ and $\mathcal{A}_{i}=\sum_{j}\mathcal{A}[i,j]$, to preserve the row-wise normalization.
@@
A hyper-parameter $\beta$ controls the degree of suppression applied to image-token weights; unless otherwise stated, we set $\beta=0.1$. After these two steps, the modified attention weight replaces the original, while all other MHA operations remain unchanged. This allows refinement of the output representation without altering the model parameters.
Figure 2: The left $5\times 5$ matrix shows the original attention weight, while the right depicts it after manipulation. CLS and IMG denote the class token and an image token, respectively, with a total length of 5. Darker blue indicates lower attention scores (approaching zero), while darker red indicates higher scores (approaching one).
@@
Table 3: Comparison on MS COCO for cross-modal retrieval. R@k denotes recall rate at rank k, and mR (mean-R) denotes the average of R@1, R@5, and R@10. Model sizes are denoted as B (ViT-B), L (ViT-L), and H (ViT-H). Best results are highlighted in bold.
4.2 AAT with Genetic Algorithm
As discussed in Section 3, ablating detrimental heads in CLIP’s image encoder can refine its output representations. Our goal is to identify the optimal subset of such heads. We formulate this as a combinatorial optimization problem and solve it using a small validation set $\mathcal{D}$ containing a limited number of image–text pairs. To efficiently explore the large search space, we employ a Genetic Algorithm (GA) [holland1992adaptation], which evolves candidate solutions over generations to discover high-quality configurations.
Specifically, we represent all attention heads in the image encoder as a binary vector, where each element corresponds to a head (1 for ablated, 0 for retained). Using the validation set $\mathcal{D}$, the GA optimizes this vector configuration. To align with CLIP’s training objective which brings matching image–text pairs closer while pushing mismatched pairs apart, we design a fitness function $\mathcal{F}$ that reflects this property. $\mathcal{F}$ serves as the optimization target and is defined as follows:
where $N$ is the number of sample pairs in $\mathcal{D}$. $S_{pos}^{i}$ denotes the cosine similarity between the $i$-th text and its ground-truth image. $S_{neg}^{i,j}$ denotes the cosine similarity between the $i$-th text and the $j$-th image in $\mathcal{H}^{i}$ (an updatable set of hard negatives for the $i$-th text), which is detailed below.
We first evaluate vanilla CLIP on $\mathcal{D}$ for text-to-image retrieval. For each text query $i$, we collect the top $k_{1}$ non-matching images (false positives) as hard negatives to initialize $\mathcal{H}^{i}$. To avoid overfitting, we augment $\mathcal{H}^{i}$ with $k_{2}$ randomly sampled non-matching images from $\mathcal{D}$ and refresh these $k_{2}$ samples at the end of each GA generation. This dynamic $\mathcal{H}$ is crucial to AAT-GA: it encourages larger margins between true matches and evolving hard negatives while continuously incorporating newly challenging cases.
In Eq. 3, the first term of $\mathcal{F}$ rewards similarity between true image–text pairs, while the second penalizes similarity to the hardest negatives in $\mathcal{H}$. Combined with standard crossover and mutation, this objective steers the GA toward an optimal (or near-optimal) head selection for ablation.
Application Scenarios
AAT-GA operates entirely at inference, making it well suited to resource-constrained deployments. It can run on the low-power, inference-oriented edge devices with limited-precision compute capabilities, such as INT8/INT16 chips (e.g., Google Coral Edge TPU, Ethos-N78, Mythic M1108 AMP). It is also data-efficient, requiring only a small validation set for optimization (Section 5.4.2), whereas supervised finetuning (SFT) tends to overfit in such settings (Appendix 12).
4.3 AAT with Back-Propagation
Since GA may require more inference trials when poorly initialized, in addition, not all resource-constrained settings lack floating-point compute, thus we introduce a gradient-based alternative: Back-Propagation (AAT-BP). Rather than using the uniform suppression factor $\beta$ in AAT-GA, AAT-BP assigns each probed head a learnable parameter $\alpha_{i}$. We map $\alpha_{i}$ to a head-specific suppression factor $\beta_{i}\in[0,1]$ that controls the ablation strength of head $i$, replacing $\beta$ with $\beta_{i}$. Formally, $\beta_{i}$ is defined as:
$\beta_{i}$ $=\text{Sigmoid}(\tau\cdot\alpha_{i})$ (4)
where the temperature $\tau>1$ sharpens the sigmoid, encouraging $\beta_{i}$ to approach extreme values (0 or 1) during optimization. Empirically, we set $\tau=5.0$ and initialize each $\alpha_{i}$ to 1.0, yielding $\beta_{i}$ near 1 at the start. As in AAT-GA, we optimize $\alpha_{i}$} on the same validation set $\mathcal{D}$ using CLIP’s contrastive supervision objective.
Application Scenarios
While BP might get stuck to local optima, its rapid convergence makes it practical and efficient in real-world use. Its low compute and data demands further suit resource-limited settings. In spirit, AAT-BP resembles PEFT, yet it uses orders of magnitude fewer learnable parameters than typical PEFT [hulora, houlsby2019parameter], a level of efficiency enabled by targeted manipulation of attention heads.